OpenStudy (anonymous):

How would you write the following expression as a sum or difference? log(3√2-x/3x) A.1/3(log(2-x)+log3x) B. 1/3log(2-x)-log3x C. 1/3log(2-x)-logx D. 1/3log(2-x)+log3x

OpenStudy (tkhunny):

Well, since you have ignored the Order of Operations and written \(3\sqrt{2} - \dfrac{x}{3}\cdot x = 3\sqrt{2} - \dfrac{x^{2}}{3}\), then none of those choices is suitable. Please give it another go, making sure you write what you intend - always recalling the Order of Operations and inserting parentheses where necessary.